you are trying to invest $36000 for 16 years and the bank informed you will receive a annual rate of return You want to buy a car in 16 years equal to $9000 LO3 a 10.32 b 5.89 x 12.45 d 7.46 e None of the above
Answer #1
In the given question, the initial investment should be invested for 16 years so that
the amount is equal to the car price i.e 9,000
Also, there seems to be a printing mistake in question wherein the initial investment
should have been 3600 instead of 36000
Amount= Principal * (1+ rate)^time
9000= 3600*(1+rate)^16
(1+rate)^16= 9000/3600
(1+rate)^16=          2.50
(1+rate)= 2.50^(1/16)
(1+rate)= 1.05894
Rate= 0.05894
Rate= 5.89%
Hence option B is correct.